# Approximation in Hölder spaces

### 03 Jan 2024 v1

@@ 3.1. Lipschitz and smooth approximations
Now, if $\|x\|\geq R+1,$ then $\|y\|\geq R$, and so $f(y)=0,$ whenever $y\in B(x,1).$ Thus, by the previous formula, we deduce $f_{n}(x)=\inf_{y\in B(x,1)}\{n\|x-y\|\}=0,$ which shows that $f_{n}$ has bounded support, for $n$ large enough. ∎
@@
By Theorem 1.5 we know that if $f\in\dot{\operatorname{VC}}^{0,\omega}(X),$ and $\varepsilon>0,$ we can find $g\in\dot{\operatorname{VC}}^{0,\omega}(X)$ with bounded support and $\|f-g\|_{\dot{C}^{0,\omega}(X)}\leq\varepsilon.$ Then, Lemma 3.1 provides us with $h\in\operatorname{Lip}_{\operatorname{bs}}(X)$ so that $\|h-g\|_{\dot{C}^{0,\omega}(X)}\leq\varepsilon,$ and therefore $\|f-h\|_{\dot{C}^{0,\omega}(X)}\leq 2\varepsilon;$ this shows the inclusion ``"."```$\supset$".` ` ⊃ " . For the converse implication ``"``"`$\subset$"` ` ⊂ ", it is enough to observe that $\mathrm{Lip}_{\operatorname{bs}}(X)\subset\dot{\operatorname{VC}}^{0,\omega}(X)$ and that $\dot{\operatorname{VC}}^{0,\omega}(X)$ is a closed subspace of $\dot{C}^{0,\omega}(X).$
@@ 3.2. Approximation in super-reflexive spaces
As we mentioned in the introduction, in Hilbert spaces the Lasry–Lions regularization theorem (see [LasryLions] or [HajJoh10, p. 408]) provides uniform approximation of Lipschitz functions by $C^{1,1}\cap\operatorname{Lip}$ functions. Cepedello-Boiso [Cepedello, Theorem 1] proved that a variant of the Lasry-Lions technique can be used to obtain uniform approximation of Lipschitz functions in super-reflexive spaces by $C^{1}$ functions whose derivatives are $\alpha$-Hölder on bounded subsets, but not globally in $X.$ These approximations are, of course, Lipschitz on bounded sets, but not globally Lipschitz. Therefore, our next goal is to approximate Lipschitz functions uniformly by globally $C^{1,\alpha}\cap\operatorname{Lip}$ functions. Moreover, we show that the Lipschitz constant is preserved up to a multiplicative factor, that depends only on the constants $\alpha$ and $C$ from (38). In order to do so, we rely on recent results on $C^{1,\omega}$ extensions for $1$-jets [AzaMud21] to construct special $C^{1,\alpha}$ separating functions with absolute bounds for their Lipschitz and $C^{1,\omega}$ seminorms. Besides the basic interest to the theory of uniform approximation, the following Theorem 3.8 will be essential for the approximation of $\dot{\operatorname{VC}}^{0,\omega}$ functions.
@@
Notice that the modulus $\omega$ is independent of $\alpha.$ Also, by Remark 3.9 the approximations in part (ii) are Lipschitz as well.
@@ 3.3. Approximation of Banach-valued mappings from $c_{0}$
For an arbitrary set of indices $\Gamma,$ the space $c_{0}(\Gamma)$ consists of those elements $x=(x_{\gamma})_{\gamma\in\Gamma}\in\mathbb{R}^{\Gamma}$ such that for every $\varepsilon>0,$ there exists a finite subset $S=S_{x}\subset\Gamma$ so that $|x_{\gamma}|\leq\varepsilon,$ whenever $\gamma\in\Gamma\setminus S.$ We equip $c_{0}(\Gamma)$ with the norm $\|(x_{\gamma})_{\gamma\in\Gamma}\|_{\infty}:=\sup_{\gamma\in\Gamma}|x_{\gamma}|$ and this results in a Banach space.
Given $x=(x_{\gamma})_{\gamma\in\Gamma}$ and $\alpha\in\Gamma,$ we denote the projection $(x)_{\alpha}=((x_{\gamma})_{\gamma\in\Gamma})_{\alpha}:=x_{\alpha}\in\mathbb{R};$ and for every finite subset $S\subset\Gamma,$ we denote
$P_{S}:c_{0}(\Gamma)\to c_{00}(\Gamma),\qquad P_{S}(x)=\sum_{\gamma\in S}x_{\gamma}e_{\gamma},$
where $e_{\gamma}$ is the element of $c_{0}(\Gamma)$ that satisfies $(e_{\gamma})_{\alpha}=\delta_{\gamma,\alpha},$ for every $\alpha\in\Gamma.$
Theorem 3.12 (Approximation in $c_{0}(\Gamma)$).
For an arbitrary set of indices $\Gamma,$ let $X=c_{0}(\Gamma),$ and let $Y$ be a Banach space. Then, the following hold:
@@
One step in our proof of Theorem 3.12 will rely on the construction by Hájek and Johanis [HajJoh09, Theorem 1] (stated for Lipschitz functions), to obtain the $C^{\infty}$ approximation on $c_{0}(\Gamma).$ However, the verification $\dot{C}^{0,\omega}$ convergence after our $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}$ condition will require more work.
@@
Let $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ and $\varepsilon>0.$ Then, there exists $g\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ and $r>0$ so that $\|g-f\|_{\dot{C}^{0,\omega}(X,Y)}\leq\varepsilon,$ and for every $x\in X,$ there exists a finite subset $S=S_{x}$ of $\Gamma$ for which $g(z)=g(P_{S}(z)),$ whenever $z\in B(x,r).$ Moreover, if $f$ has bounded support (resp. Lipschitz), then the above $g$ has bounded supported (resp. Lipschitz) too.
@@
$\phi:X\to X,\qquad\phi(x)=\phi((x_{\gamma})_{\gamma\in\Gamma})=(\varphi(x_{\gamma}))_{\gamma\in\Gamma}.$
@@
We have shown that $\|f-g\|_{\dot{C}^{0,\omega}(X,Y)}\leq\varepsilon.$ Now, if $x\in X=c_{0}(\Gamma),$ then (by definition) there exists a finite subset $S=S_{x}$ such that if $\gamma\in\Gamma\setminus S,$ then $|x_{\gamma}|\leq r/2.$ Then, for $z\in B(x,r/2)$ and for each $\gamma\in\Gamma\setminus S$ it follows that $|z_{\gamma}|\leq r,$ implying that $\varphi(z_{\gamma})=0.$ It follows for $z\in B(x,r/2)$ that $\phi(z)=\phi(P_{S}(z))$ and hence that $g(z)=g(P_{S}(z)).$ For the second part, suppose there exists $R>0$ so that $f(x)=0$ for all $\|x\|\geq R.$ If $\|z\|\geq R+r,$ then $|z_{\gamma_{0}}|\geq R+r$ for some $\gamma_{0}\in\Gamma.$ Consequently
$\|\phi(z)\|=\sup_{\gamma\in\Gamma}|\varphi(z_{\gamma})|\geq|\varphi(z_{\gamma_{0}})|\geq|z_{\gamma_{0}}|-r\geq R,$
@@
In the proof of part (i), by Lemma 3.13, given a fixed $r>0,$ it is enough to approximate functions $g\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ with the property that for every $x\in X,$ there exists finite $S_{x}\subset\Gamma$ such that $g(z)=g(P_{S_{x}}(z)),$ for all $z\in B(x,2r).$ Given $\varepsilon,$ let $\delta>0$ be so that
@@
Moreover, for $0<\eta<\min(r,\varepsilon)$ small enough (more precisely, so that $\omega(\eta)\leq\varepsilon\omega(\delta)/(1+\|g\|_{\dot{C}^{0,\omega}})),$ let $\theta:\mathbb{R}\to\mathbb{R}$ be an even $C^{\infty}$ smooth non-negative function such that $\int_{\mathbb{R}}\theta=1$ and $\theta=0$ on $\mathbb{R}\setminus[-\eta,\eta].$ For every finite set $F\subset\Gamma,$ and $x\in X,$ we define the Bochner integral
$h_{F}(x):=\int_{[-\eta,\eta]^{|F|}}g\Big{(}x-\sum_{\gamma\in F}t_{\gamma}e_{\gamma}\Big{)}\prod_{\gamma\in F}\theta(t_{\gamma})\,\mathrm{d}\lambda_{|F|}(t),$ (45)
@@
$h_{S_{x}}(z)=\int_{[-\eta,\eta]^{|S_{x}|}}g\Big{(}P_{S_{x}}(z)-\sum_{\gamma\in S_{x}}t_{\gamma}e_{\gamma}\Big{)}\prod_{\gamma\in S_{x}}\theta(t_{\gamma})\,\mathrm{d}\lambda_{|S_{x}|}(t),\quad z\in B(x,r).$
Since $g\in\dot{C}^{0,\omega}(X,Y)$ and $P_{S_{x}}$ is $1$-Lipschitz, the function $z\mapsto(g\circ P_{S_{x}})(z)$ belongs to $\dot{C}^{0,\omega}(X,Y)$, implying that $z\mapsto(g\circ P_{S_{x}})(z)$ is uniformly continuous on $X.$ Therefore, by the previous formula, $B(x,r)\ni z\mapsto h_{S_{x}}(z)$ is a finite dimensional smooth mollification (in the Bochner sense) with a uniformly continuous mapping, which shows that $h_{S_{x}}\in C^{\infty}(B(x,r),Y).$ Moreover, using Fubini’s theorem, one can easily check that if $F\subset\Gamma$ is finite and $F\supset S_{x},$ then also $h_{F}=h_{S_{x}}$ on $B(x,r).$ This enables us to define the desired mapping $h:X\to Y$ in the following manner. Given $x\in X,$ let $S_{x}\subset\Gamma$ be the subset described above, and define $h(x):=h_{S_{x}}(x).$ Moreover, ordering the collection of finite subsets $\mathcal{F}(\Gamma)$ of $\Gamma$ by inclusion, the pointwise limit of the net $\{h_{S}\}_{S\in\mathcal{F}(\Gamma)}$ is exactly $h.$
@@
Since $z\mapsto h_{S_{x}}(z)$ is of class $C^{\infty}(B(x,r),Y)$, the above yields $h\in C^{\infty}(X,Y).$ All these properties were stated and proved in [HajJoh09, Lemma 6]. We next show that $h\in\dot{\operatorname{VC}}_{\operatorname{small}}^{0,\omega}(X,Y).$ Given any two points $x,z\in X,$ let $S_{x},S_{z}\subset\Gamma$ be the finite subsets associated with $x$ and $z,$ and set $S=S_{x}\cup S_{z}.$ Since $h(x)=h_{S}(x)$ and $h(z)=h_{S}(z)$ we have
$\begin{split}\|h(x)-h(z)\|_{Y}&=\Big{\|}\int_{[-\eta,\eta]^{|S|}}\Big{(}g\big{(}x-\sum_{\gamma\in S}t_{\gamma}e_{\gamma}\big{)}-g\big{(}z-\sum_{\gamma\in S}t_{\gamma}e_{\gamma}\big{)}\Big{)}\prod_{\gamma\in S}\theta(t_{\gamma})\,\mathrm{d}\lambda_{|S|}(t)\Big{\|}_{Y}\\ &\leq\int_{[-\eta,\eta]^{|S|}}\Big{\|}g\big{(}x-\sum_{\gamma\in S}t_{\gamma}e_{\gamma}\big{)}-g\big{(}z-\sum_{\gamma\in S}t_{\gamma}e_{\gamma}\big{)}\Big{\|}_{Y}\prod_{\gamma\in S}\theta(t_{\gamma})\,\mathrm{d}\lambda_{|S|}(t)\\ &\leq\sup_{\begin{subarray}{c}u,v\in X,\\ \|u-v\|=\|x-z\|\end{subarray}}\|g(u)-g(v)\|_{Y},\end{split}$ (46)
bearing in mind that $\int_{[-\eta,\eta]^{|S|}}\prod_{\gamma\in S}\theta(t_{\gamma})\,\mathrm{d}\lambda_{|S|}(t)=1.$ It follows immediately from the bound (46) that $\|h\|_{\dot{C}^{0,\omega}(X,Y)}\leq\|g\|_{\dot{C}^{0,\omega}(X,Y)}$ and $h\in\dot{\operatorname{VC}}_{\operatorname{small}}^{0,\omega}(X,Y).$ Now if $\|x-z\|\leq\delta,$ then together (46) and (44) give
@@
$\|h(x)-g(x)\|_{Y}$ $\leq\int_{[-\eta,\eta]^{|S|}}\Big{\|}g\Big{(}x-\sum_{\gamma\in S}t_{\gamma}e_{\gamma}\Big{)}-g(x)\Big{\|}_{Y}\prod_{\gamma\in S}\theta(t_{\gamma})\,\mathrm{d}\lambda_{|S|}(t)$
$\leq\|g\|_{\dot{C}^{0,\omega}(X,Y)}\int_{[-\eta,\eta]^{|S|}}\omega\Big{(}\Big{\|}\sum_{\gamma\in S}t_{\gamma}e_{\gamma}\Big{\|}\Big{)}\prod_{\gamma\in S}\theta(t_{\gamma})\,\mathrm{d}\lambda_{|S|}(t)$
@@
We conclude that $\|g-h\|_{\dot{C}^{0,\omega}(X,Y)}\leq 2\varepsilon$ and complete the proof of Theorem 3.12(i). For part (ii), given $f\in\dot{\operatorname{VC}}^{0,\omega}(X,Y),$ by Theorem 1.5, we approximate $f$, in the $\dot{C}^{0,\omega}(X,Y)$-seminorm by a $g\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ with bounded support. By Lemma 3.13, we can assume that for every $x\in X$ there exists a finite $S=S_{x}$ such that $g=g\circ P_{S}$ on $B(x,r),$ and for a uniform fixed $r>0.$ Let us next check that the approximation $h$ from part (i) has bounded support. Let $R>0$ be so that $g$ is zero on $X\setminus B(0,R).$ Assume that $\|x\|\geq R+\eta.$ There exists a finite set $S=S_{x}\subset\Gamma$ so that $h(x)=h_{S}(x)$. Now, for $t=(t_{\gamma})_{\gamma\in S}\in[-\eta,\eta]^{|S|},$ there holds that
$\Big{\|}x-\sum_{\gamma\in S}t_{\gamma}e_{\gamma}\Big{\|}\geq\|x\|-\Big{\|}\sum_{\gamma\in S}t_{\gamma}e_{\gamma}\Big{\|}\geq\|x\|-\eta\geq R$
@@
$h(x)=h_{S}(x)=\int_{[-\eta,\eta]^{|S|}}g\Big{(}x-\sum_{\gamma\in S}t_{\gamma}e_{\gamma}\Big{)}\prod_{\gamma\in S}\theta(t_{\gamma})\,\mathrm{d}\lambda_{|S|}(t)=0.$
@@
For the particular case of $\Gamma=\{1,\ldots,n\},$ $c_{0}(\Gamma)$ becomes $\mathbb{R}^{n}$ with the supremum norm, which is of course equivalent with the usual Euclidean setting. Theorem 3.12 then has the following consequence.

### 21 Feb 2024 v2

@@ 3.1. Lipschitz and smooth approximations
Now, if $\|x\|\geq R+1,$ then $\|y\|\geq R$, and so $f(y)=0,$ whenever $y\in B(x,1).$ Thus, by the previous formula, we deduce $f_{n}(x)=\inf_{y\in B(x,1)}\{n\|x-y\|\}=0,$ which shows that $f_{n}$ has bounded support, for $n$ large enough.
@@
By Theorem 1.4 we know that if $f\in\dot{\operatorname{VC}}^{0,\omega}(X),$ and $\varepsilon>0,$ we can find $g\in\dot{\operatorname{VC}}^{0,\omega}(X)$ with bounded support and $\|f-g\|_{\dot{C}^{0,\omega}(X)}\leq\varepsilon.$ Then, Lemma 3.1 provides us with $h\in\operatorname{Lip}_{\operatorname{bs}}(X)$ so that $\|h-g\|_{\dot{C}^{0,\omega}(X)}\leq\varepsilon,$ and therefore $\|f-h\|_{\dot{C}^{0,\omega}(X)}\leq 2\varepsilon;$ this shows the inclusion ``"."```$\supset$".` ` ⊃ " .
For the converse implication ``"``"`$\subset$"` ` ⊂ ", it is enough to observe that $\mathrm{Lip}_{\operatorname{bs}}(X)\subset\dot{\operatorname{VC}}^{0,\omega}(X)$ and that $\dot{\operatorname{VC}}^{0,\omega}(X)$ is a closed subspace of $\dot{C}^{0,\omega}(X).$
@@ 3.2. Approximation in super-reflexive spaces
In Hilbert spaces the Lasry–Lions regularization theorem (see [LasryLions] or [HajJoh10, p. 408]) provides uniform approximation of Lipschitz functions by $C^{1,1}\cap\operatorname{Lip}$ functions. Cepedello-Boiso [Cepedello, Theorem 1] used an ingenious variant of the Lasry-Lions technique to obtain uniform approximation of Lipschitz functions in super-reflexive spaces by $C^{1}$ functions whose derivatives are $\alpha$-Hölder on bounded subsets, but not globally on $X.$ These approximations are, of course, Lipschitz on bounded sets, but not globally Lipschitz. However, for our purposes, we need to approximate Lipschitz functions uniformly by globally $C^{1,\alpha}\cap\operatorname{Lip}$ functions, and we need a good control on the Lipschitz constants of the approximations. In Theorem 3.8(2) below, we have obtained uniform approximations that are globally $C^{1,\alpha}$ and globally Lipschitz on $X$, preserving the Lipschitz constants up to a multiplicative factor depending only on $\alpha$ and $X.$ In the proof, we use a recent result on $C^{1,\omega}$ extensions for jets [AzaMud21] to construct $C^{1,\alpha}$ bump functions with all the properties stated in part (1) of Theorem 3.8. This is combined with a method of Hájek and Johanis [HajJoh10, Proposition 1] to glue up a suitable sequence of these bumps. Both parts (1) and (2) in Theorem 3.8 are interesting to the theory of smooth approximation, and also will be essential for the approximation of $\dot{\operatorname{VC}}^{0,\omega}$ functions.
Let us mention that, after making the first version of this article public, it has come to our attention that part (2) of Theorem 3.8 was very recently discovered by Johanis [Jo24] too, by means of a different proof. A benefit of Johanis’ proof is that it provides the sharp constant $\kappa=1.$ Since we only need an absolute control on $\kappa$ for our purposes, and since we consider part (1) of independent interest as well, we have chosen to include our original proof of Theorem 3.8(2).
@@
Notice that the modulus $\omega$ is independent of $\alpha.$ Also, the approximations in part (ii) are Lipschitz. Indeed, as the derivative is $\alpha$-Hölder and boundedly supported, it is in particular bounded and this implies that the function is Lipschitz.
@@ 3.3. Approximation of Banach-valued mappings from $c_{0}$
For an arbitrary set of indices $\mathcal{A},$ the space $c_{0}(\mathcal{A})$ consists of those elements $x=(x_{\alpha})_{\alpha\in\mathcal{A}}\in\mathbb{R}^{\mathcal{A}}$ such that for every $\varepsilon>0,$ there exists a finite subset $S=S_{x}\subset\mathcal{A}$ so that $|x_{\alpha}|\leq\varepsilon,$ whenever $\alpha\in\mathcal{A}\setminus S.$ We equip $c_{0}(\mathcal{A})$ with the norm $\|(x_{\alpha})_{\alpha\in\mathcal{A}}\|_{\infty}:=\sup_{\alpha\in\mathcal{A}}|x_{\alpha}|$ and this results in a Banach space.
Given $x=(x_{\alpha})_{\alpha\in\mathcal{A}}$ and $\alpha\in\mathcal{A},$ we denote the projection $(x)_{\beta}=((x_{\alpha})_{\alpha\in\mathcal{A}})_{\beta}:=x_{\beta}\in\mathbb{R};$ and for every finite subset $S\subset\mathcal{A},$ we denote
$P_{S}:c_{0}(\mathcal{A})\to c_{00}(\mathcal{A}),\qquad P_{S}(x)=\sum_{\alpha\in S}x_{\alpha}e_{\alpha},$
where $e_{\alpha}$ is the element of $c_{0}(\mathcal{A})$ that satisfies $(e_{\alpha})_{\beta}=\delta_{\alpha,\beta},$ for every $\beta\in\mathcal{A}.$
Theorem 3.12 (Approximation in $c_{0}(\mathcal{A})$).
For an arbitrary set of indices $\mathcal{A},$ let $X=c_{0}(\mathcal{A}),$ and let $Y$ be a Banach space. Then, the following hold:
@@
One step in our proof of Theorem 3.12 will rely on the construction by Hájek and Johanis [HajJoh09, Theorem 1] (stated for Lipschitz functions), to obtain the $C^{\infty}$ approximation on $c_{0}(\mathcal{A}).$ However, the verification $\dot{C}^{0,\omega}$ convergence after our $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}$ condition will require more work.
@@
Let $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ and $\varepsilon>0.$ Then, there exists $g\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ and $r>0$ so that $\|g-f\|_{\dot{C}^{0,\omega}(X,Y)}\leq\varepsilon,$ and for every $x\in X,$ there exists a finite subset $S=S_{x}$ of $\mathcal{A}$ for which $g(z)=g(P_{S}(z)),$ whenever $z\in B(x,r).$ Moreover, if $f$ has bounded support (resp. Lipschitz), then the above $g$ has bounded supported (resp. Lipschitz) too.
@@
$\phi:X\to X,\qquad\phi(x)=\phi((x_{\alpha})_{\alpha\in\mathcal{A}})=(\varphi(x_{\alpha}))_{\alpha\in\mathcal{A}}.$
@@
We have shown that $\|f-g\|_{\dot{C}^{0,\omega}(X,Y)}\leq\varepsilon.$ Now, if $x\in X=c_{0}(\mathcal{A}),$ then (by definition) there exists a finite subset $S=S_{x}$ such that if $\alpha\in\mathcal{A}\setminus S,$ then $|x_{\alpha}|\leq r/2.$ Then, for $z\in B(x,r/2)$ and for each $\alpha\in\mathcal{A}\setminus S$ it follows that $|z_{\alpha}|\leq r,$ implying that $\varphi(z_{\alpha})=0.$ It follows for $z\in B(x,r/2)$ that $\phi(z)=\phi(P_{S}(z))$ and hence that $g(z)=g(P_{S}(z)).$ For the second part, suppose there exists $R>0$ so that $f(x)=0$ for all $\|x\|\geq R.$ If $\|z\|\geq R+r,$ then $|z_{\alpha_{0}}|\geq R+r$ for some $\alpha_{0}\in\mathcal{A}.$ Consequently
$\|\phi(z)\|=\sup_{\alpha\in\mathcal{A}}|\varphi(z_{\alpha})|\geq|\varphi(z_{\alpha_{0}})|\geq|z_{\alpha_{0}}|-r\geq R,$
@@
In the proof of part (i), by Lemma 3.13, given a fixed $r>0,$ it is enough to approximate functions $g\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ with the property that for every $x\in X,$ there exists finite $S_{x}\subset\mathcal{A}$ such that $g(z)=g(P_{S_{x}}(z)),$ for all $z\in B(x,2r).$ Given $\varepsilon,$ let $\delta>0$ be so that
@@
Moreover, for $0<\eta<\min(r,\varepsilon)$ small enough (more precisely, so that $\omega(\eta)\leq\varepsilon\omega(\delta)/(1+\|g\|_{\dot{C}^{0,\omega}})),$ let $\theta:\mathbb{R}\to\mathbb{R}$ be an even $C^{\infty}$ smooth non-negative function such that $\int_{\mathbb{R}}\theta=1$ and $\theta=0$ on $\mathbb{R}\setminus[-\eta,\eta].$ For every finite set $F\subset\mathcal{A},$ and $x\in X,$ we define the Bochner integral
$h_{F}(x):=\int_{[-\eta,\eta]^{|F|}}g\Big{(}x-\sum_{\alpha\in F}t_{\alpha}e_{\alpha}\Big{)}\prod_{\alpha\in F}\theta(t_{\alpha})\,\mathrm{d}\lambda_{|F|}(t),$ (46)
@@
$h_{S_{x}}(z)=\int_{[-\eta,\eta]^{|S_{x}|}}g\Big{(}P_{S_{x}}(z)-\sum_{\alpha\in S_{x}}t_{\alpha}e_{\alpha}\Big{)}\prod_{\alpha\in S_{x}}\theta(t_{\alpha})\,\mathrm{d}\lambda_{|S_{x}|}(t),\quad z\in B(x,r).$
Since $g\in\dot{C}^{0,\omega}(X,Y)$ and $P_{S_{x}}$ is $1$-Lipschitz, the function $z\mapsto(g\circ P_{S_{x}})(z)$ belongs to $\dot{C}^{0,\omega}(X,Y)$, implying that $z\mapsto(g\circ P_{S_{x}})(z)$ is uniformly continuous on $X.$ Therefore, by the previous formula, $B(x,r)\ni z\mapsto h_{S_{x}}(z)$ is a finite dimensional smooth mollification (in the Bochner sense) with a uniformly continuous mapping, which shows that $h_{S_{x}}\in C^{\infty}(B(x,r),Y).$ Moreover, using Fubini’s theorem, one can easily check that if $F\subset\mathcal{A}$ is finite and $F\supset S_{x},$ then also $h_{F}=h_{S_{x}}$ on $B(x,r).$ This enables us to define the desired mapping $h:X\to Y$ in the following manner. Given $x\in X,$ let $S_{x}\subset\mathcal{A}$ be the subset described above, and define $h(x):=h_{S_{x}}(x).$ Moreover, ordering the collection of finite subsets $\mathcal{F}(\mathcal{A})$ of $\mathcal{A}$ by inclusion, the pointwise limit of the net $\{h_{S}\}_{S\in\mathcal{F}(\mathcal{A})}$ is exactly $h.$
@@
Since $z\mapsto h_{S_{x}}(z)$ is of class $C^{\infty}(B(x,r),Y)$, the above yields $h\in C^{\infty}(X,Y).$ All these properties were stated and proved in [HajJoh09, Lemma 6]. We next show that $h\in\dot{\operatorname{VC}}_{\operatorname{small}}^{0,\omega}(X,Y).$ Given any two points $x,z\in X,$ let $S_{x},S_{z}\subset\mathcal{A}$ be the finite subsets associated with $x$ and $z,$ and set $S=S_{x}\cup S_{z}.$ Since $h(x)=h_{S}(x)$ and $h(z)=h_{S}(z)$ we have
$\begin{split}\|h(x)-h(z)\|_{Y}&=\Big{\|}\int_{[-\eta,\eta]^{|S|}}\Big{(}g\big{(}x-\sum_{\alpha\in S}t_{\alpha}e_{\alpha}\big{)}-g\big{(}z-\sum_{\alpha\in S}t_{\alpha}e_{\alpha}\big{)}\Big{)}\prod_{\alpha\in S}\theta(t_{\alpha})\,\mathrm{d}\lambda_{|S|}(t)\Big{\|}_{Y}\\ &\leq\int_{[-\eta,\eta]^{|S|}}\Big{\|}g\big{(}x-\sum_{\alpha\in S}t_{\alpha}e_{\alpha}\big{)}-g\big{(}z-\sum_{\alpha\in S}t_{\alpha}e_{\alpha}\big{)}\Big{\|}_{Y}\prod_{\alpha\in S}\theta(t_{\alpha})\,\mathrm{d}\lambda_{|S|}(t)\\ &\leq\sup_{\begin{subarray}{c}u,v\in X,\\ \|u-v\|=\|x-z\|\end{subarray}}\|g(u)-g(v)\|_{Y},\end{split}$ (47)
bearing in mind that $\int_{[-\eta,\eta]^{|S|}}\prod_{\alpha\in S}\theta(t_{\alpha})\,\mathrm{d}\lambda_{|S|}(t)=1.$ It follows immediately from the bound (47) that $\|h\|_{\dot{C}^{0,\omega}(X,Y)}\leq\|g\|_{\dot{C}^{0,\omega}(X,Y)}$ and $h\in\dot{\operatorname{VC}}_{\operatorname{small}}^{0,\omega}(X,Y).$ Now if $\|x-z\|\leq\delta,$ then together (47) and (45) give
@@
$\|h(x)-g(x)\|_{Y}$ $\leq\int_{[-\eta,\eta]^{|S|}}\Big{\|}g\Big{(}x-\sum_{\alpha\in S}t_{\alpha}e_{\alpha}\Big{)}-g(x)\Big{\|}_{Y}\prod_{\alpha\in S}\theta(t_{\alpha})\,\mathrm{d}\lambda_{|S|}(t)$
$\leq\|g\|_{\dot{C}^{0,\omega}(X,Y)}\int_{[-\eta,\eta]^{|S|}}\omega\Big{(}\Big{\|}\sum_{\alpha\in S}t_{\alpha}e_{\alpha}\Big{\|}\Big{)}\prod_{\alpha\in S}\theta(t_{\alpha})\,\mathrm{d}\lambda_{|S|}(t)$
@@
We conclude that $\|g-h\|_{\dot{C}^{0,\omega}(X,Y)}\leq 2\varepsilon$ and complete the proof of Theorem 3.12(i). For part (ii), given $f\in\dot{\operatorname{VC}}^{0,\omega}(X,Y),$ by Theorem 1.4, we approximate $f$, in the $\dot{C}^{0,\omega}(X,Y)$-seminorm by a $g\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ with bounded support. By Lemma 3.13, we can assume that for every $x\in X$ there exists a finite $S=S_{x}$ such that $g=g\circ P_{S}$ on $B(x,r),$ and for a uniform fixed $r>0.$ Let us next check that the approximation $h$ from part (i) has bounded support. Let $R>0$ be so that $g$ is zero on $X\setminus B(0,R).$ Assume that $\|x\|\geq R+\eta.$ There exists a finite set $S=S_{x}\subset\mathcal{A}$ so that $h(x)=h_{S}(x)$. Now, for $t=(t_{\alpha})_{\alpha\in S}\in[-\eta,\eta]^{|S|},$ there holds that
$\Big{\|}x-\sum_{\alpha\in S}t_{\alpha}e_{\alpha}\Big{\|}\geq\|x\|-\Big{\|}\sum_{\alpha\in S}t_{\alpha}e_{\alpha}\Big{\|}\geq\|x\|-\eta\geq R$
@@
$h(x)=h_{S}(x)=\int_{[-\eta,\eta]^{|S|}}g\Big{(}x-\sum_{\alpha\in S}t_{\alpha}e_{\alpha}\Big{)}\prod_{\alpha\in S}\theta(t_{\alpha})\,\mathrm{d}\lambda_{|S|}(t)=0.$
@@
For the particular case of $\mathcal{A}=\{1,\ldots,n\},$ $c_{0}(\mathcal{A})$ becomes $\mathbb{R}^{n}$ with the supremum norm, which is of course equivalent with the usual Euclidean setting. Theorem 3.12 then has the following consequence.